The Everton College store paid $1966 for an order of 52 calculators. The store paid $8 for each scientific calculator. The others, all graphing calculators, cost the store $58 each. How many of each type of calculator was ordered? The store order _____ scientific calculators and _____ graphing calculators.
Answer by stanbon(75887)

The Everton College store paid $1966 for an order of 52 calculators. The store paid $8 for each scientific calculator. The others, all graphing calculators, cost the store $58 each. How many of each type of calculator was ordered? The store order _____ scientific calculators and _____ graphing calculators.
------------------
Let # of scientific be "s".
Let # of graphing be "g".
--------------------------
EQUATIONS:
Quantity Eq.: s + g = 52
Value Eq.: 8s + 58g = 1966
-----------------
Rearrange:
s = 52-g
Substitute to solve for "g":
8(52-g) + 58g = 1966
416 - 8g + 58g = 1966
50g = 1550
g = 31 (# of graphing calculators)
----------
Substitute into s = 52-g to solve for "s":
52-g = 52-31 = 21 (# of scientific calculators)
=====================
Cheers,
Stan H.